OpenStudy (thomas5267):

Solve the following linear recurrence relations. \[ A_n=-x A_{n-1}-\frac{1}{4}A_{n-2} \] This is a recurrence relations of the characteristic polynomial of a particular kind of matrix. EDIT: I need to prove that \(M_n\) are always diagonalisable. I am only interested in \(n\geq3\text{ and odd}\). \(M_n\) is a tridiagonal nxn matrix. \(M_n\) has 1/2 on all entries on the upper off-diagonal except the last entry, which is equal to 1. It also has 1/2 on all entries on the lower off-diagonal except the first entry, which is also equal to 1. All other entries are zero. For example: \[ M_5=\begin{pmatrix} 0&\frac{1}{2}&0&0&0\\ 1&0&\frac{1}{2}&0&0\\ 0&\frac{1}{2}&0&\frac{1}{2}&0\\ 0&0&\frac{1}{2}&0&1\\ 0&0&0&\frac{1}{2}&0 \end{pmatrix}\\ M_7=\begin{pmatrix} 0&\frac{1}{2}&0&0&0&0&0\\ 1&0&\frac{1}{2}&0&0&0&0\\ 0&\frac{1}{2}&0&\frac{1}{2}&0&0&0\\ 0&0&\frac{1}{2}&0&\frac{1}{2}&0&0\\ 0&0&0&\frac{1}{2}&0&\frac{1}{2}&0\\ 0&0&0&0&\frac{1}{2}&0&1\\ 0&0&0&0&0&\frac{1}{2}&0 \end{pmatrix}\\ \] The characteristic equation of \(M_n\) is \(-xA_{n-1}-\frac{1}{2}A_{n-2}, A_n=-x A_{n-1}-\frac{1}{4}A_{n-2}\). \(A_n\) is the characteristic polynomial of the submatrix of \(M_n\). The submatrix is generated by dropping the first row and first column of \(M_n\). For example: \[ M_5=\begin{pmatrix} 0&\frac{1}{2}&0&0&0\\ 1&0&\frac{1}{2}&0&0\\ 0&\frac{1}{2}&0&\frac{1}{2}&0\\ 0&0&\frac{1}{2}&0&1\\ 0&0&0&\frac{1}{2}&0 \end{pmatrix}\\ M_5-xI_5=\begin{pmatrix} -x&\frac{1}{2}&0&0&0\\ 1&-x&\frac{1}{2}&0&0\\ 0&\frac{1}{2}&-x&\frac{1}{2}&0\\ 0&0&\frac{1}{2}&-x&1\\ 0&0&0&\frac{1}{2}&-x \end{pmatrix}\\ \begin{align*} \det(M_5-xI_5)&= -x\begin{vmatrix} -x&\frac{1}{2}&0&0\\ \frac{1}{2}&-x&\frac{1}{2}&0\\ 0&\frac{1}{2}&-x&1\\ 0&0&\frac{1}{2}&-x \end{vmatrix}-\frac{1}{2} \begin{vmatrix} 1&\frac{1}{2}&0&0\\ 0&-x&\frac{1}{2}&0\\ 0&\frac{1}{2}&-x&1\\ 0&0&\frac{1}{2}&-x \end{vmatrix}\\ &=-xA_4-\frac{1}{2}\begin{vmatrix} -x&\frac{1}{2}&0\\ \frac{1}{2}&-x&1\\ 0&\frac{1}{2}&-x\\ \end{vmatrix}\\ &=-xA_4-\frac{1}{2}A_3 \end{align*} \] \(A_n\) has a close form solution of \(A_n = \left( \dfrac{-x+\sqrt{x^2-1}}{2}\right)^n + \left(\dfrac{-x-\sqrt{x^2-1}}{2}\right)^n\) with the initial condition \(A_1=-x,\,A_2=x^2-\frac{1}{2}\).

OpenStudy (thomas5267):

x is not related to n.

ganeshie8 (ganeshie8):

characteristic eqn is \(r^2+xr+\frac{1}{4} = 0\) \(\implies r = \dfrac{-x\pm\sqrt{x^2-1}}{2}\)

ganeshie8 (ganeshie8):

therefore the general solution of recurrence relation is \[A_n = c_1\left( \dfrac{-x+\sqrt{x^2-1}}{2}\right)^n + c_2\left(\dfrac{-x-\sqrt{x^2-1}}{2}\right)^n\]

OpenStudy (thomas5267):

Okay. The initial condition is \(A_1=-x,\,A_2=x^2-\frac{1}{2}\). I plugged \(n=1\) and \(x=1\) and \(x=\sqrt{2}\) to get \(c_1=c_2=1\).

OpenStudy (thomas5267):

Is it possible to simplify \(A_n\)? Any identity relating to \((a+b)^n-(a-b)^n\)?

OpenStudy (thomas5267):

I mean \((a+b)^n+(a-b)^n\).

ganeshie8 (ganeshie8):

it looks good the way it is now i don't see any easy way to simplify it further recall the formula for fibonacci sequence

ganeshie8 (ganeshie8):



OpenStudy (thomas5267):

Any idea how to prove this polynomial has n distinct root?

OpenStudy (thomas5267):

I want to prove that matrix is always diagonalisable.

ganeshie8 (ganeshie8):

oh if you show that eigenvalues are distinct, then that essentially shows that enough eigenvectors are available

ganeshie8 (ganeshie8):

Is \(n\) the size of matrix here ?

OpenStudy (thomas5267):

Yes n is the size of the matrix.

ganeshie8 (ganeshie8):

and you want to prove that the polynomial eqn \(A_n=0\) always has \(n\) distinct roots ?

OpenStudy (thomas5267):

I want to prove these matrices are always diagonalisable. \[ M_5=\begin{pmatrix} 0&\frac{1}{2}&0&0&0\\ 1&0&\frac{1}{2}&0&0\\ 0&\frac{1}{2}&0&\frac{1}{2}&0\\ 0&0&\frac{1}{2}&0&1\\ 0&0&0&\frac{1}{2}&0 \end{pmatrix}\\ M_7=\begin{pmatrix} 0&\frac{1}{2}&0&0&0&0&0\\ 1&0&\frac{1}{2}&0&0&0&0\\ 0&\frac{1}{2}&0&\frac{1}{2}&0&0&0\\ 0&0&\frac{1}{2}&0&\frac{1}{2}&0&0\\ 0&0&0&\frac{1}{2}&0&\frac{1}{2}&0\\ 0&0&0&0&\frac{1}{2}&0&1\\ 0&0&0&0&0&\frac{1}{2}&0 \end{pmatrix}\\ M_7-xI_7=\begin{pmatrix} -x&\frac{1}{2}&0&0&0&0&0\\ 1&-x&\frac{1}{2}&0&0&0&0\\ 0&\frac{1}{2}&-x&\frac{1}{2}&0&0&0\\ 0&0&\frac{1}{2}&-x&\frac{1}{2}&0&0\\ 0&0&0&\frac{1}{2}&-x&\frac{1}{2}&0\\ 0&0&0&0&\frac{1}{2}&-x&1\\ 0&0&0&0&0&\frac{1}{2}&-x \end{pmatrix}\\ \] \(A_n\) is the first entry of the cofactor expansion of \(M_n-xI_n\). \[ A_6=\begin{vmatrix} -x&\frac{1}{2}&0&0&0&0\\ \frac{1}{2}&-x&\frac{1}{2}&0&0&0\\ 0&\frac{1}{2}&-x&\frac{1}{2}&0&0\\ 0&0&\frac{1}{2}&-x&\frac{1}{2}&0\\ 0&0&0&\frac{1}{2}&-x&1\\ 0&0&0&0&\frac{1}{2}&-x \end{vmatrix}\\ \]

OpenStudy (thomas5267):

Proving \(A_n\) always has n distinct solution would be a good start I think.

ganeshie8 (ganeshie8):

Ahh these are special matrices they are called tridiagonal matrices or something, very interesting ones..

OpenStudy (thomas5267):

It is a transition matrix of a Markov chain. I used this matrix to model the gunplay of a first person shooter called Planetside 2 XD.

ganeshie8 (ganeshie8):

Wow! cool stuff xD you found that recurrence relation by working 1x1, 2x2, 3x3, and then using induction is it ?

OpenStudy (thomas5267):

I didn't bother with induction. I expanded the matrix corresponded to \(A_4\) and found the empirical relations.

ganeshie8 (ganeshie8):

Ohk, I remember a nice simple way to get that earlier reccurrence relation let me pull that up quick

ganeshie8 (ganeshie8):

here it is https://en.wikipedia.org/wiki/Tridiagonal_matrix#Determinant

ganeshie8 (ganeshie8):

okay lets get back to proving that \(A_n=0\) has \(n\) distinct roots

OpenStudy (thomas5267):

I am only interested in the diagonalisability of \(M_n,\,n\geq3 \text{ and is odd}\). Actually proving \(A_n\) has n distinct roots is not that useful since \(\det(M_n-xI_n)\) is not \(A_n\).

ganeshie8 (ganeshie8):

this looks interesting https://en.wikipedia.org/wiki/Tridiagonal_matrix#Eigenvalues

ganeshie8 (ganeshie8):

I thought \(A_n\) is the characteristic polynomial of nxn matrix ?

ganeshie8 (ganeshie8):

by that, i mean solutions of \(A_n=0\) are the eigenvalues that you're interested in

ganeshie8 (ganeshie8):

or did i get that whole thing wrong ?

OpenStudy (thomas5267):

\[ \det(M-xI_n)=-xA_{n-1}-\frac{1}{4}A_{n-2} \]

OpenStudy (thomas5267):

I am only interested in the diagonalisability of \(M_n, n\geq3\text{ and is odd}\). \(A_n\) is the characteristic polynomial of the submatrix of \(M_{n+1}\) with the first row and first column removed.

OpenStudy (thomas5267):

Sorry I didn't made it clear.

OpenStudy (thomas5267):

Actually the characteristic polynomial of \(M_n\) is not what I stated above.

OpenStudy (thomas5267):

The relationship should be \(\det(M_n-xI_n)=-xA_{n-1}-\frac{1}{2}A_{n-2}\)

ganeshie8 (ganeshie8):

Okay I think I'm beginning to understand.. looks more involved than I thought..

OpenStudy (thomas5267):

Just to reiterate. \[ A_n = \left( \dfrac{-x+\sqrt{x^2-1}}{2}\right)^n + \left(\dfrac{-x-\sqrt{x^2-1}}{2}\right)^n\]

ganeshie8 (ganeshie8):

not so sure how you got that relation but it seems you want to show \(-xA_{n-1}-\frac{1}{2}A_{n-2}=0\) has \(n\) distinct solutions

OpenStudy (thomas5267):

Yes. I want to show \(\det(M_n-xI_n)=-xA_{n-1}-\frac{1}{2}A_{n-2}=0\) has n distinct solution.

ganeshie8 (ganeshie8):

something is wrong

ganeshie8 (ganeshie8):

nvm

OpenStudy (thomas5267):

\[ M_5=\begin{pmatrix} 0&\frac{1}{2}&0&0&0\\ 1&0&\frac{1}{2}&0&0\\ 0&\frac{1}{2}&0&\frac{1}{2}&0\\ 0&0&\frac{1}{2}&0&1\\ 0&0&0&\frac{1}{2}&0 \end{pmatrix}\\ M_5-xI_5=\begin{pmatrix} -x&\frac{1}{2}&0&0&0\\ 1&-x&\frac{1}{2}&0&0\\ 0&\frac{1}{2}&-x&\frac{1}{2}&0\\ 0&0&\frac{1}{2}&-x&1\\ 0&0&0&\frac{1}{2}&-x \end{pmatrix}\\ \begin{align*} \det(M_5-xI_5)&= -x\begin{vmatrix} -x&\frac{1}{2}&0&0\\ \frac{1}{2}&-x&\frac{1}{2}&0\\ 0&\frac{1}{2}&-x&1\\ 0&0&\frac{1}{2}&-x \end{vmatrix}-\frac{1}{2} \begin{vmatrix} 1&\frac{1}{2}&0&0\\ 0&-x&\frac{1}{2}&0\\ 0&\frac{1}{2}&-x&1\\ 0&0&\frac{1}{2}&-x \end{vmatrix}\\ &=-xA_4-\frac{1}{2}\begin{vmatrix} -x&\frac{1}{2}&0\\ \frac{1}{2}&-x&1\\ 0&\frac{1}{2}&-x\\ \end{vmatrix}\\ &=-xA_4-\frac{1}{2}A_3 \end{align*} \\ \]

OpenStudy (thomas5267):

\(A_n\) are all polynomials from 1 to 10.

ganeshie8 (ganeshie8):

right, i see..

ganeshie8 (ganeshie8):

thats really a clever derivation for determinant ! il give it a try in the evening again, right now i must go..

OpenStudy (thomas5267):

Thank you! It is not really that hard to find once you tried to find the characteristic polynomial of \(M_n\) by induction.

OpenStudy (thomas5267):

\[ \begin{align*} \det(M_n-x I_n)&=-xA_{n-1}-\frac{1}{2}A_{n-2}\\ &=A_n-\frac{1}{4}A_{n-2} \end{align*} \]

OpenStudy (thomas5267):

The eigenvalues of \(M_{n},\,n\geq3 \text{ and odd}\) are symmetric (x and -x are both eigenvalues.) and distinct. 0 is always one of the eigenvalues. The characteristic polynomial is also odd.

OpenStudy (thomas5267):

@Kainui Help!